### v1

# NLP-ADBench: NLP Anomaly Detection Benchmark

Abstract.
Anomaly detection (AD) is a critical machine learning task with diverse applications in web systems, including fraud detection, content moderation, and user behavior analysis. Despite its significance, AD in natural language processing (NLP) remains underexplored, limiting advancements in detecting anomalies in text data such as harmful content, phishing attempts, or spam reviews. In this paper, we introduce NLP-ADBench, the most comprehensive benchmark for NLP anomaly detection (NLP-AD), comprising eight curated datasets and evaluations of nineteen state-of-the-art algorithms. These include three end-to-end methods and sixteen two-step algorithms that apply traditional anomaly detection techniques to language embeddings generated by bert-base-uncased and OpenAI’s text-embedding-3-large models.
Our results reveal critical insights and future directions for NLP-AD. Notably, no single model excels across all datasets, highlighting the need for automated model selection. Moreover, two-step methods leveraging transformer-based embeddings consistently outperform specialized end-to-end approaches, with OpenAI embeddings demonstrating superior performance over BERT embeddings. By releasing NLP-ADBench at https://github.com/USC-FORTIS/NLP-ADBench, we provide a standardized framework for evaluating NLP-AD methods, fostering the development of innovative approaches. Th is work fills a crucial gap in the field and establishes a foundation for advancing NLP anomaly detection, particularly in the context of improving the safety and reliability of web-based systems.
Anomaly Detection, Natural Language Processing, Web Mining, NLP Benchmark, Benchmark, Datasets
*Equal contribution.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Anomaly detectionccs: Information systems Web miningccs: Computing methodologies Natural language processing
1. Introduction
Anomaly detection (AD) is a fundamental area in machine learning with diverse applications in web systems, such as fraud detection, content moderation, and user behavior analysis (Chandola et al., 2009; Ahmed et al., 2016). While significant progress has been made in AD for structured data types, including tabular, graph, and time series data (Chalapathy and Chawla, 2019; Han et al., 2022; Lai et al., 2021; Liu et al., 2022), its application to natural language processing (NLP) tasks—a domain crucial to many web-based applications—remains both essential and relatively under explored (Ruff et al., 2021). For instance, detecting harmful or abusive language is critical for ensuring safe user experiences in social media platforms and online forums (Fortuna and Nunes, 2018). Similarly, in e-commerce, identifying anomalies in reviews or product descriptions can help detect spam or fraudulent behaviors, critical for preserving user trust and platform integrity.
Key Contributions. In this work, we introduce NLP-ADBench, a comprehensive benchmark for NLP anomaly detection (NLP-AD). Building on our prior benchmarking efforts in tabular (Han et al., 2022), graph (Liu et al., 2022), and time-series (Lai et al., 2021) anomaly detection, NLP-ADBench (as the most comprehensive NLP-AD benchmark) provides several key advantages over existing works (Bejan et al., 2023): (i) a diverse collection of 8 datasets reflecting real-world web scenarios, (ii) integration of 19 advanced NLP-based anomaly detection methods, (iii) actionable insights for advancing NLP-AD research, and (iv) fully open-sourced datasets, algorithm implementation, and experiments.
Key Insights/Takeaways (see details in §3). Our experiments reveal the following key findings: (i) There is no single model that consistently performs best across all datasets due to variability in dataset characteristics, highlighting the need for automated model selection. (ii) Transformer-based embeddings significantly enhance the performance of two-step AD methods, such as LUNAR (Goodge et al., 2022) and LOF (Breunig et al., 2000), compared to end-to-end approaches. (iii) High-dimensional embeddings, such as those from OpenAI models, improve detection accuracy but require balancing performance with computational efficiency. (iv) Addressing dataset-specific biases and reducing human dependency in anomaly definitions remain critical for advancing robust and generalizable NLP-AD systems.
Our study aligns closely with the Web Mining and Content Analysis” track, particularly addressing topics such as Robustness and generalizability of Web mining methods” and “Web data quality in the era of algorithmically-generated content.”
2. NLP-ADBench: AD Benchmark for NLP Tasks
2.1. Preliminaries and Problem Definition
Anomaly Detection in Natural Language Processing (NLP-AD) focuses on identifying text instances that deviate significantly from expected or typical patterns. Unlike structured data, text data is inherently unstructured, high-dimensional, and deeply influenced by the nuances of human language, including syntax, semantics, and context (Aggarwal, 2017).These unique properties introduce significant challenges, making the development of robust and accurate AD methods for NLP a complex and demanding task.
Formally, let $\mathcal{D}={x_{1},x_{2},\dots,x_{N}}$ denote a corpus where each $x_{i}$ is a text instance. The goal of NLP-AD is to learn an anomaly scoring function $f:\mathcal{X}\rightarrow\mathbb{R}$ that assigns a real-valued anomaly score to each text instance. Higher scores correspond to greater deviations from normal patterns, indicating a higher likelihood of an instance being anomalous.
Table 1. Statistical information of the NLPAD dataset.
NLPAD Dataset # Samples #Normal #Anomaly %Anomaly
NLPAD-AGNews $98,207$ $94,427$ $3,780$ $3.85\%$
NLPAD-BBCNews $1,785$ $1,723$ $62$ $3.47\%$
NLPAD-EmailSpam $3,578$ $3,432$ $146$ $4.08\%$
NLPAD-Emotion $361,980$ $350,166$ $11,814$ $3.26\%$
NLPAD-MovieReview $26,369$ $24,882$ $1,487$ $5.64\%$
NLPAD-N24News $59,822$ $57,994$ $1,828$ $3.06\%$
NLPAD-SMSSpam $4,672$ $4,518$ $154$ $3.30\%$
NLPAD-YelpReview $316,924$ $298,986$ $17,938$ $5.66\%$
2.2. Curated Benchmark Datasets
The limited availability of purpose-built datasets constrains the development and evaluation of effective methods in NLP-AD. To address this gap, we curated and transformed 8 existing classification datasets from various NLP domains into specialized datasets tailored for NLP-AD tasks, ensuring that all data are presented in a standard format. These datasets, collectively called the NLPAD datasets, provide a foundational resource for advancing research.
Each transformed dataset is named by adding the prefix “NLPAD-” to the original dataset’s name (e.g., NLPAD-AGNews, NLPAD-BBCNews), distinguishing them from the original datasets. The NLPAD datasets are provided in a unified JSON Lines format [^1] A format for storing structured data, see https://jsonlines.org/. for compatibility and ease of use. Each line is a JSON object [^2] A lightweight data format, see https://www.json.org/json-en.html. with four fields: text (the text used for anomaly detection), label (the anomaly detection label, where 1 represents an anomaly and 0 represents normal), original_task (the task of the original dataset), and original_label (the category label from the original dataset).
To transform each dataset for NLP-AD, we established a text selection process based on the data format. For tabular data, we carefully chose appropriate columns as the text source. For document-based data, we extracted text directly from relevant documents. The anomalous class for each dataset was selected based on semantic distinctions within the dataset categories, ensuring that the identified anomalies represent meaningful deviations from the normal data distribution (Emmott et al., 2015; Han et al., 2022). Once identified, the anomalous class was downsampled to represent less than 10% of the total instances.
For instance, in the NLPAD-AGNews dataset (tabular data), we selected the “description” column as the text source, with the “World” category serving as the anomalous class due to its semantic divergence from other categories such as “Sports” or “Technology.” This anomalous class was then downsampled accordingly. Similarly, in the NLPAD-BBCNews dataset (document-based data), text from BBC News documents was used, with the “entertainment” category identified as anomalous because its semantic content significantly differs from other categories like “Politics” or “Business.” The “entertainment” category was also downsampled to maintain consistency. This semantic-driven approach to defining anomalies was consistently applied across all datasets. Further details of dataset sources and construction processes can be found in Appx. A.1.1. Table. 1 presents the statistical information of the NLPAD datasets, including the total number of samples, the number of normal and anomalous samples, and the anomaly ratio for each dataset.
2.3. The Most Comprehensive NLP-AD Algorithms with Open Implementations
Table 2. Performance comparison of 19 Algorithms on 8 NLPAD datasets using AUROC, with best results highlighted in bold and shaded.
Methods NLPAD-AGNews NLPAD-BBCNews NLPAD-EmailSpam NLPAD-Emotion NLPAD-MovieReview NLPAD-N24News NLPAD-SMSSpam NLPAD-YelpReview
CVDD 0.6046 0.7221 0.9340 0.4867 0.4895 0.7507 0.4782 0.5345
DATE 0.8120 0.9030 0.9697 0.6291 0.5185 0.7493 0.9398 0.6092
FATE 0.7756 0.9310 0.9061 0.5035 0.5289 0.8073 0.6262 0.5945
BERT + LOF 0.7432 0.9320 0.7482 0.5435 0.4959 0.6703 0.7190 0.6573
BERT + DeepSVDD 0.6671 0.5683 0.6937 0.5142 0.4287 0.4366 0.5859 0.5871
BERT + ECOD 0.6318 0.6912 0.7052 0.5889 0.4282 0.4969 0.5606 0.6326
BERT + iForest 0.6124 0.6847 0.6779 0.4944 0.4420 0.4724 0.5053 0.5971
BERT + SO-GAAL 0.4489 0.3099 0.4440 0.5031 0.4663 0.4135 0.3328 0.4712
BERT + AE 0.7200 0.8839 0.4739 0.5594 0.4650 0.5749 0.6918 0.6441
BERT + VAE 0.6773 0.7409 0.4737 0.5594 0.4398 0.4949 0.6082 0.6441
BERT + LUNAR 0.7694 0.9260 0.8417 0.5186 0.4687 0.6284 0.6953 0.6522
OpenAI + LOF 0.8905 0.9558 0.9263 0.7304 0.6156 0.7806 0.7862 0.8733
OpenAI + DeepSVDD 0.4680 0.5766 0.4415 0.4816 0.6563 0.6150 0.3491 0.5373
OpenAI + ECOD 0.7638 0.7224 0.9263 0.6206 0.7366 0.7342 0.4317 0.5984
OpenAI + iForest 0.5213 0.6064 0.6937 0.5889 0.5064 0.4944 0.3751 0.5871
OpenAI + SO-GAAL 0.5945 0.2359 0.4440 0.5031 0.6201 0.5043 0.5671 0.5082
OpenAI + AE 0.8326 0.9520 0.7651 0.7067 0.6088 0.7155 0.5511 0.8524
OpenAI + VAE 0.8144 0.7250 0.5273 0.7067 0.4515 0.7418 0.4259 0.6163
OpenAI + LUNAR 0.9226 0.9732 0.9343 0.9328 0.6474 0.8320 0.7189 0.9452
Compared to the existing NLP-AD benchmark by Bejan et al. (Bejan et al., 2023), NLP-ADBench provides a broader evaluation by including 19 algorithms, categorized into two groups. The first group comprises 3 end-to-end algorithms that directly process raw text data to produce anomaly detection outcomes. The second group consists of 16 algorithms derived by applying 8 traditional anomaly detection (AD) methods to text embeddings generated from two models: bert-base-uncased (Devlin et al., 2019) and OpenAI’s text-embedding-3-large (OpenAI, 2024). These traditional AD methods do not operate on raw text directly but instead perform anomaly detection on embeddings, offering a complementary approach to the end-to-end methods. This comprehensive algorithm collection enables a robust evaluation of direct and embedding-based NLP anomaly detection techniques. Here, we provide a brief description; see details in Appx. A.2.
End-to-end NLP-AD Algorithms. We evaluate three end-to-end algorithms tailored for NLP anomaly detection. (1) Context Vector Data Description (CVDD) (Ruff et al., 2019) leverages context vectors and pre-trained embeddings with a multi-head self-attention mechanism to project normal instances close to learned contexts, identifying anomalies based on deviations. (2) Detecting Anomalies in Text via Self-Supervision of Transformers (DATE) (Manolache et al., 2021) trains transformers using self-supervised tasks like replaced mask detection to capture normal text patterns and flag anomalies. (3) Few-shot Anomaly Detection in Text with Deviation Learning (FATE) (Das et al., 2023) uses a few labeled anomalies with deviation learning to distinguish anomalies from normal instances, achieving state-of-the-art performance on benchmark datasets.
Two-step NLP-AD Algorithms. We evaluate 8 two-step algorithms that rely on embeddings generated by models such as bert-base-uncased (Devlin et al., 2019) and text-embedding-3-large (OpenAI, 2024). These algorithms are designed to work with structured numerical data and cannot directly process raw textual data, requiring text transformation into numerical embeddings. (4) LOF (Breunig et al., 2000) measures local density deviations, while (5) DeepSVDD (Ruff et al., 2018) minimizes the volume of a hypersphere enclosing normal representations. (6) ECOD (Li et al., 2022) uses empirical cumulative distribution functions to estimate densities and assumes anomalies lie in distribution tails. (7) IForest (Liu et al., 2008) recursively isolates anomalies through random splits, and (8) SO_GAAL (Liu et al., 2019) generates adversarial samples to identify anomalies. Reconstruction-based approaches include (9) AE (Aggarwal, 2017), which flags anomalies based on reconstruction errors, and (10) VAE (Kingma and Welling, 2013; Burgess et al., 2018), which identifies anomalies using reconstruction probabilities or latent deviations. Finally, (11) LUNAR (Goodge et al., 2022) enhances traditional local outlier detection with graph neural networks.
Accessibility. We open-source NLP-ADBench with datasets and algorithm implementation for reproducibility and future research.
3. Experiment Results
Figure 1. Average rank on AUCROC of 19 NLPAD methods across 8 datasets (the lower the better).
3.1. Experiment Setting
Datasets, Train/Test Data Split, and Independent Trials. In the NLP-ADBench benchmark, the data is divided by allocating 70% of the normal data to the training set. The remaining 30% of normal data, combined with all anomalous data, forms the test set. To ensure the robustness of our findings, we repeat each experiment three times and report the average performance.
Hyperparameter Settings. For all the algorithms in NLP-ADBench, we use their default hyperparameter (HP) settings in the original paper for a fair comparison, same as ADBench (Han et al., 2022).
Evaluation Metrics and Statistical Tests. We evaluate different NLP-AD methods by a widely used metric: AUCROC (Area Under Receiver Operating Characteristic Curve).
Embeddings Definitions:
  1. (1)
    BERT refers specifically to the bert-base-uncased model (Devlin et al., 2019).
  2. (2)
    OpenAI refers to OpenAI’s text-embedding-3-large model (OpenAI, 2024).
  3. (3)
    The term “BERT + AD algorithm” or “OpenAI + AD algorithm” means that we first generate text embeddings using BERT or OpenAI’s model, respectively, and then apply the AD algorithm.
3.2. Results, Discussions, and New Directions
No single model performs best across all datasets due to variability in dataset characteristics. The performance of AD models varies significantly depending on dataset characteristics, making it challenging to identify a universal winner. For datasets with a larger variety of categories, such as NLPAD-AGNews, OpenAI + LUNAR achieved a score of 0.9226, outperforming the end-to-end method CVDD (0.6046) by 52.6%. Similarly, on NLPAD-BBCNews, OpenAI + LOF scored 0.9558, exceeding CVDD (0.7221) by 32.4%. Conversely, for datasets with fewer categories or binary-class structures, end-to-end methods show distinct advantages. For example, on the binary-class NLPAD-SMSSpam dataset, the end-to-end method DATE achieved 0.9398, outperforming OpenAI + LUNAR (0.7189) by 30.7%.
  • Future Direction 1: Automated Model Selection. These results emphasize the importance of developing automated approaches to select the most suitable model based on the characteristics of the dataset, e.g., the number of categories and the complexity of the structure. One feasible solution will be adapting the meta-learning framework from tabular AD settings (Zhao et al., 2021) to NLPAD.
Transformer-based embeddings boost the performance of two-step AD methods. Two-step AD algorithms, when paired with transformer-based embeddings, have shown the ability to outperform end-to-end methods like CVDD and FATE in NLP AD tasks. For instance, OpenAI + LUNAR achieved a significantly higher performance of 0.9226 on the NLPAD-AGNews dataset, outperforming CVDD by 52.6% and FATE by 19.0%. Similarly, OpenAI + LOF achieved 0.9558 on the NLPAD-BBCNews dataset, surpassing CVDD by 32.4% and FATE by 2.7%. These results emphasize the substantial advantage of leveraging transformer-based embeddings over relying solely on end-to-end methods. The success of two-step approaches can be attributed to the high-quality contextual embeddings produced by transformer models like OpenAI, which serve as a robust foundation for AD. In contrast, some end-to-end methods, such as CVDD, rely on embeddings from older models like GloVe, highlighting the need for incorporating more advanced embeddings to improve performance.
  • Future Direction 2: Efficient and Robust Embedding Utilization. While transformer-based embeddings deliver superior accuracy, their computational cost and resource requirements remain significant. Future research should focus on designing lightweight, efficient algorithms that can harness updated embeddings without excessive time or resource consumption. Additionally, improving the robustness of two-step approaches to ensure consistent performance across diverse datasets and anomaly types will be crucial for advancing NLP anomaly detection systems.
High-dimensional embeddings improve detection but require balancing performance and efficiency. The dimensionality of embeddings varies significantly across models, impacting both performance and efficiency in AD. While the BERT-base model generates 768-dimensional embeddings, OpenAI’s text-embedding-3-large produces embeddings with a dimensionality of 3072—a 300% increase. Our experiments reveal that OpenAI embeddings consistently outperform BERT embeddings across multiple datasets in NLP-ADBench, demonstrating the advantages of higher-dimensional embeddings. For instance, OpenAI + LUNAR achieved a performance of 0.9452 on the NLPAD-YelpReview dataset, surpassing BERT + LUNAR’s 0.6522 by 44.9%. Similarly, on NLPAD-AGNews, OpenAI + LUNAR scored 0.9226, exceeding BERT + LUNAR’s 0.7694 by 19.9%. On NLPAD-N24News, OpenAI + LUNAR reached 0.8320, outperforming BERT + LUNAR’s 0.6284 by 32.4%. These results highlight the capability of high-dimensional embeddings to enhance AD performance. However, the increased dimensionality also introduces challenges, such as higher computational costs and potential information redundancy.
  • Future Direction 3: Optimizing Embedding Dimensionality. Future research should explore dimensionality reduction techniques tailored to NLP-AD tasks to mitigate redundancy and computational overhead without compromising performance. Additionally, hybrid approaches that adaptively adjust dimensionality based on dataset characteristics could enable scalable, efficient, and effective AD systems.
4. Conclusion, Limitations, and Future Works
We present NLP-ADBench, the most comprehensive benchmark for contextual NLP anomaly detection (NLP-AD), evaluating 19 state-of-the-art algorithms across 8 diverse datasets. Our findings establish the superiority of two-step methods leveraging transformer-based embeddings, such as OpenAI + LUNAR, over end-to-end approaches, demonstrating the power of hybrid strategies for handling complex NLP anomaly detection tasks. By combining advanced text embeddings with traditional anomaly detection methods, NLP-ADBench provides a robust and flexible framework that sets a new standard for evaluating NLP-AD systems. Additionally, we offer actionable insights into model performance, dataset variability, and embedding utilization, paving the way for future research.
Despite its contributions, NLP-ADBench has certain limitations. First, the datasets included in the benchmark, while diverse, are primarily sourced from existing classification tasks and may not fully reflect emerging challenges such as anomalies in multilingual or multimodal text data. Second, our evaluations focus on static embeddings, leaving dynamic or streaming NLP-AD scenarios unexplored. Third, the reliance on predefined anomaly labels in our benchmark limits the ability to assess unsupervised or domain-adaptive approaches. Future work can expand NLP-ADBench to include more diverse datasets, such as multilingual or multimodal data, and by exploring dynamic anomaly detection in streaming text scenarios. Incorporating benchmarks for unsupervised and adaptive models can also better reflect real-world applications. These advancements will enhance NLP-ADBench’s utility as a comprehensive platform for driving progress in NLP anomaly detection.
Supplementary Material for NLP-ADBench
Appendix A Details on NLP-ADBench
A.1. Additional Details on Benchmark Datasets
A.1.1. Datasets Sources.
  1. (1)
    NLPAD-AGNews is constructed from the AG News dataset (Rai, 2023), which was originally intended for news topic classification tasks. The AG News dataset contains 127,600 samples categorized into four classes: World, Sports, Business, and Sci/Tech. We selected the text from the “description” column as NLPAD-AGNews’s text data source. The “World” category was designated as the anomaly class and was downsampled accordingly.
  2. (2)
    NLPAD-BBCNews is constructed from the BBC News dataset (Greene and Cunningham, 2006), which was originally used for document classification across various news topics. The BBC News dataset includes 2,225 articles divided into five categories: Business, Entertainment, Politics, Sport, and Tech. We selected the full text of the news articles as NLPAD-BBC News’s text data source. The “Entertainment” category was designated as the anomaly class and was downsampled accordingly.
  3. (3)
    NLPAD-EmailSpam is constructed from the Spam Emails dataset (Metsis et al., 2006), originally used for email spam detection. The Spam Emails dataset contains 5,171 emails labeled as either spam or ham (not spam). We selected the text from the “text” column containing the email bodies as NLPAD-Emails Spam’s text data source. The “spam” category was designated as the anomaly class and was downsampled accordingly.
  4. (4)
    NLPAD-Emotion: is constructed from the Emotion dataset (Saravia et al., 2018) , which was originally intended for emotion classification tasks in textual data. The Emotion dataset contains 416,809 text samples labeled with six emotions: anger, fear, joy, love, sadness, and surprise. We selected the text from the “text” column as NLPAD-Emotion’s text data source. The “fear” category was designated as the anomaly class and was downsampled accordingly.
  5. (5)
    NLPAD-MovieReview: is constructed from the Movie Review dataset (Maas et al., 2011) , commonly used for sentiment analysis of film critiques. The Movie Review dataset includes 50,000 reviews labeled as positive or negative. We selected the full review texts as NLPAD-MovieReview’s text data source. The “neg” (negative reviews) category was designated as the anomaly class and was downsampled accordingly.
  6. (6)
    NLPAD-N24News is constructed from the N24News dataset (Wang et al., 2022), originally used for topic classification of news articles. N24News contains 61,235 articles across various categories. We selected the full text of the news articles as NLPAD-N24News’s text data source. The “food” category was designated as the anomaly class and was downsampled accordingly.
  7. (7)
    NLPAD-SMSSpam is constructed from the SMS Spam Collection dataset (Almeida et al., 2011), originally intended for classifying SMS messages as spam or ham (not spam). The SMS Spam Collection dataset comprises 5,574 messages labeled accordingly. We selected the text from the “message text” as NLPAD-SMS Spam’s text data source. The “spam” category was designated as the anomaly class and was downsampled accordingly.
  8. (8)
    NLPAD-YelpReview is constructed from the Yelp Review Polarity dataset  (Putra, 2023), originally intended for sentiment classification tasks. The Yelp Review Polarity dataset is created by considering 1-star and 2-star ratings as negative, and 3-star and 4-star ratings as positive. For each polarity, 280,000 training samples and 19,000 testing samples were randomly selected, resulting in a total of 560,000 training samples and 38,000 testing samples. Negative polarity is labeled as class 1, and positive polarity as class 2. We selected the text from the text column as NLPAD-YelpReview’s text data source. The label 1 (negative reviews) was designated as the anomaly class and was downsampled accordingly.
A.1.2. NLPAD dataset’s text pre-processing
On all 8 datasets, we preprocessed the raw text data to ensure consistency and usability by removing URLs and HTML tags, eliminating unnecessary special characters while retaining essential punctuation, converting line breaks and consecutive spaces into single spaces, and preserving case sensitivity and stop words to maintain linguistic integrity. After processing the text, we found that some texts became duplicates due to the removal of certain symbols. Consequently, we removed all duplicate data to ensure the uniqueness of each text sample. These preprocessing steps follow established practices to effectively clean text data while retaining its syntactic and semantic features, providing a reliable foundation for natural language processing tasks (Chai, 2022).
A.2. Additional Details on Algorithms
A.2.1. End-to-End Algorithms
  1. (1)
    Context Vector Data Description (Ruff et al., 2019)(CVDD) is an unsupervised anomaly detection method for textual data. It utilizes pre-trained word embeddings and a multi-head self-attention mechanism to learn ”context vectors” that represent normal patterns in the data. Anomalies are detected by measuring the cosine distance between sequence projections and context vectors, where larger distances indicate higher anomaly likelihoods. CVDD also penalizes overlapping contexts to enhance interpretability.
  2. (2)
    Detecting Anomalies in Text via Self-Supervision of Transformers (DATE) (Manolache et al., 2021) detects anomalies in text by training self-supervised transformers on tasks like replaced mask detection, enabling the model to learn normal language patterns and identify deviations.
  3. (3)
    Few-shot Anomaly Detection in Text with Deviation Learning (FATE) (Das et al., 2023) is a deep learning framework that uses a small number of labeled anomalies to learn anomaly scores end-to-end. By employing deviation learning, it ensures normal examples align with reference scores while anomalies deviate significantly. Utilizing multi-head self-attention and multiple instance learning, FATE achieves state-of-the-art performance on benchmark datasets.
A.2.2. Traditional Algorithms
  1. (1)
    Local Outlier Factor (LOF) (Breunig et al., 2000) calculates the local density deviation of a data point relative to its neighbors. This metric identifies points that have substantially lower density than their neighbors, marking them as outliers.
  2. (2)
    Deep Support Vector Data Description (DeepSVDD) (Ruff et al., 2018) minimizes the volume of a hypersphere enclosing the data representations learned by a neural network, capturing common patterns while identifying anomalies as points outside the hypersphere.
  3. (3)
    Empirical-Cumulative-distribution-based Outlier Detection (ECOD)  (Li et al., 2022) estimates the empirical cumulative distribution function (ECDF) for each feature independently. It identifies outliers as data points that reside in the tails of these distributions. This approach is hyperparameter-free and offers straightforward interpretability.
  4. (4)
    Isolation Forest (IForest) (Liu et al., 2008) detects anomalies by isolating observations through random feature selection and splitting, with anomalies requiring fewer splits
  5. (5)
    Single-Objective Generative Adversarial Active Learning (SO_GAAL) (Liu et al., 2019) optimizes a single objective function to generate adversarial samples and effectively identify anomalies in unsupervised settings.
  6. (6)
    AutoEncoder (AE) (Aggarwal, 2017) detects anomalies by reconstructing input data, where higher reconstruction errors signify potential anomalies.
  7. (7)
    Unifying Local Outlier Detection Methods via Graph Neural Networks(LUNAR) (Goodge et al., 2022) uses graph neural networks to integrate and enhance traditional local outlier detection methods, unifying them for better anomaly detection.
  8. (8)
    Variational AutoEncoder (VAE) (Kingma and Welling, 2013; Burgess et al., 2018) uses probabilistic latent variables to model data distributions, identifying anomalies based on reconstruction probabilities or latent space deviations.